OpenStudy (anonymous):

Sasha is planning to put a patio around her new pool which has the dimensions 20m and 30m. She wants the patio to have a uniform width, x, and to surround the entire pool. Her budget is $6000 and she knows it will cost $10/m^2 to construct the patio. How wide can she afford the patio to be?

OpenStudy (anonymous):

PLEEEASE HELP ive been trying for an hour

OpenStudy (pfenn1):

You can solve this several ways. One way is to assume the area of the patio and pool is \[(20 + 2x) \times (30+2x)\]

OpenStudy (anonymous):

oooo my god i love you

OpenStudy (pfenn1):

Subtract from this the area of just the pool = \[20 \times 30\] This will give you the area of just the patio=\[(20+2x) \times (30+3x)-(20 \times 30)\]

OpenStudy (pfenn1):

What ever you get for the area of the patio times $10/m^2 has to be equal or less than $6000. Solve for x.

OpenStudy (anonymous):

yaaay :) i love you

OpenStudy (pfenn1):

I hope it works out.